# LLM-based In-situ Thought Exchanges for Critical Paper Reading

### v1

@@ 7.1.1. Incorporating Agent(s) Feedback Promotes Critical Thinking (RQ1)
Although we did not observe clear quantitative differences in effectiveness between single and multi condition, our open-ended question analysis results uncovered participants’ perception of learning analysis skills in multi condition. As shown in Table 4, 8 out of 15 participants in multi condition reported improvements in their Analysis abilities, four times more than those in single condition. Analysis is a core higher-order cognitive process in Bloom’s Taxonomy (Bloom et al., 1956), involving the ability to break down information and compare viewpoints, which is essential for critical thinking. A possible explanation is that engaging with multiple agents in parallel and using our proposed interface to add reactions such as pin, thumb up, or thumb down (Figure 2-C2, C3) required participants to compare and analyze different comments. This feature illustrates a unique advantage of multi-agent thought exchange interfaces, which can prompt users to engage in comparative and analytical reasoning.
@@ 7.2. Limitations and perspective for future work
First, we only simulated agents representing different research domains with LLM to generate a diverse set of thoughts. Simulated AI agents are not necessarily fully equivalent to human experts in terms of knowledge and thinking capacity. In addition, the responses generated by these AI agents may overlap because the controls by LLM prompts are not strict. Future work should investigate a more algorithmic approach to generate a set of different thoughts.
Second, the present user study focused on a relatively small and specific sample of participants. Future studies could include a larger and more diverse participant pool to examine whether the patterns we observed hold across different levels of expertise or reading strategies.
@@ Acknowledgements.
We thank Zefan Sramek, Shitao Fang, Jo Takezawa, and the other members of IIS Lab for their support during this work. This work was partly supported by JST ASPIRE for Top Scientists (Grant Number JPMJAP2405), JST PRESTO (Grant Number JPMJPR23IB), JST SPRING (Grant Number JPMJSP2108) and the JSPS Invitational Fellowship for Research in Japan program (Long-term).
@@ Appendix D Prompts
Note that none, single, and multi condition share the same prompt for critical thinking questions generation. In single condition, the prompt to generate single-disciplinary feedback differs slightly from that in multi condition: it provides feedback from one disciplinary perspective that is different from the participant’s major, and without {pinned agent} input.
The text inside {} represents the participants’ input.
Prompt for Critical Thinking Questions Generation Input Section: {text} Category: {category} Output - Question Requirement 1. Read the given section content carefully. 2. Identify the main topic and the core ideas. 3. Generate only one critical thinking question that is directly relevant to the section and aligned with the assigned category. 4. Ensure that the question follows the structure and tone of the example questions provided. 5. Return only the question with no extra formatting or explanation. 6. Ensure the question is: - Clear and easy to understand for junior researchers, using simple and precise language. - Logically structured and specific, avoiding vague or overly abstract expressions. - Thought-provoking and rigorous, encouraging the reader to reflect critically. Examples Input Section: {text} Category: {category} Output {few_shot_questions}
Prompt for Multi-Disciplinary Agent Feedback The instructor asks a critical thinking question:{question}, and the reader’s answer is:{answer}. Judge the reader’s answer first, then provide an improved response from experts in three distinct and relevant academic disciplines. First, generate answers from {pinned agent} perspectives. Each answer should emphasize how it aligns with or challenges current trends or theories in that discipline. Ensure that each discipline’s answer introduces fundamentally different insights rather than paraphrasing the same idea in different terminology. Each answer should: Begin with an implicit evaluation of the reader’s answer before transitioning into the improved response. Highlight alignment with or challenges to current trends or theories in that discipline. Ensure each discipline’s answer introduces fundamentally different insights, avoiding rephrasing the same idea. Be concise (less than 100 words per answer) In the output, replace any instance of ’The reader’s answer’ with ’Your answer’. Provide the output as raw JSON text, without any surrounding Markdown formatting like triple back-ticks. Do not split or abbreviate any disciplines. Each discipline in {pinned agent} should be treated as a distinct entity. If fewer than three disciplines are provided, choose an additional relevant discipline to complete the set of three: { [Agent Name]: [Judgment of answer][Improved answer] [Agent Name]: [Judgment of answer][Improved answer] [Agent Name]: [Judgment of answer][Improved answer] }
Prompt for Reinterpreting Highlights via Multi-Disciplinary Agents The reader highlights {quote} in the paper and marks it as {label}. Reinterpret {label} of the reader’s highlight from 3 different academic disciplines’ perspective. First, generate reinterpretation from {pinned agent} perspectives. Each reinterpretation should emphasize how it aligns with or challenges current trends or theories in that discipline. Ensure that each discipline’s reinterpretation introduces fundamentally different insights rather than paraphrasing the same idea in different terminology. Each reinterpretation should: Highlight alignment with or challenges to current trends or theories in that discipline. Ensure each discipline’s reinterpretation introduces fundamentally different insights, avoiding rephrasing the same idea. Be concise (less than 100 words per reinterpretation) Provide the output as raw JSON text, without any surrounding Markdown formatting like triple back-ticks. Do not split or abbreviate any disciplines. Each discipline in {pinned agent} should be treated as a distinct entity. If fewer than three disciplines are provided, choose an additional relevant discipline to complete the set of three: { [Agent Name]: [Reinterpretation] [Agent Name]: [Reinterpretation] [Agent Name]: [Reinterpretation] }
Prompt for Answering Comments via Multi-Disciplinary Agents The reader highlights {quote} in the paper and adds comment {comment}. If the reader asks a question, answer the question based on highlights from 3 different academic disciplines’ perspective. If the reader leaves a statement, discuss their highlight and comment from 3 different academic disciplines’ perspective. First, generate a response from {pinned agent} perspectives. Each response should emphasize how it aligns with or challenges current trends or theories in that discipline. Ensure that each discipline’s response introduces fundamentally different insights rather than paraphrasing the same idea in different terminology. Each response should: Highlight alignment with or challenges to current trends or theories in that discipline. Ensure each discipline’s response introduces fundamentally different insights, avoiding rephrasing the same idea. Be concise (less than 100 words per response) Provide the output as raw JSON text, without any surrounding Markdown formatting like triple back-ticks. Do not split or abbreviate any disciplines. Each discipline in {pinned agent} should be treated as a distinct entity. If fewer than three disciplines are provided, choose an additional relevant discipline to complete the set of three: { [Agent Name]: [Response] [Agent Name]: [Response] [Agent Name]: [Response] }
@@
Prompt for Revised Answer Feedback User’s initial answer is {initialAnswer}, the revised answer is {revisedAnswer}. The feedback for the initial answer is {agreeText}. Based on this feedback, evaluate how well the revised answer addresses the issues or suggestions mentioned. Provide a concise and constructive feedback on the revised answer in 100 words.

### v2

@@ 6.3. Perceived Critical Thinking (RQ1)
6.4. Perceived Workload (RQ1)
6.5. Interaction Behavior Analysis (RQ2)
6.6. Open-Ended Question Analysis (RQ2)
@@ 7.1.1. Incorporating Agent(s) Feedback Promotes Critical Thinking (RQ1)
Although we did not observe clear quantitative differences in effectiveness between single and multi condition, our open-ended question analysis results uncovered participants’ perception of learning analysis skills in multi condition. As shown in Table 4, 8 out of 15 participants in multi condition reported improvements in their Analysis abilities, four times more than those in single condition. Analysis is a core higher-order cognitive process in Bloom’s Taxonomy (Bloom et al., 1956), involving the ability to break down information and compare viewpoints, which is essential for critical thinking. A possible explanation is that engaging with multiple agents in parallel and using our proposed interface to add reactions such as pin, thumb up, or thumb down (Figure 2-C2, C3) required participants to compare and analyze different comments. This feature illustrates a unique advantage of multi-agent thought exchange interfaces, which can prompt users to engage in comparative and analytical reasoning. From the theoretical perspective, this behavior can be interpreted as a form of metacognition scaffolding (Flavell, 1979), in which external prompts encourage readers to monitor and compare their own reasoning during reading process.
@@ 7.2. Limitations and perspective for future work
First, we only simulated agents representing different research domains with LLM to generate a diverse set of thoughts. Simulated AI agents are not necessarily fully equivalent to human experts in terms of knowledge and thinking capacity. In addition, the responses generated by these AI agents may overlap because the controls by LLM prompts are not strict. Relatedly, our AI agent design remains limited in that agents operate largely in parallel and independently, rather than engaging in an interactive scholarly dialogue. Future work should investigate more algorithmic and interaction-driven designs to generate diverse thoughts and allow agents to respond to and build on one another. Moreover, the agent mediated conditions differed not only in agent count but also in domain assignment, the multi condition used dynamically selected disciplines, whereas the single condition relied on a fixed domain outside participants’ expertise, which introduces a potential confound. Future work could address this by ensuring that participants across different conditions encounter domains matched in familiarity, difficulty, and number, in order to better understand how variations in agent expertise contribute to participants’ critical thinking outcomes.
Second, the present user study focused on a relatively small and specific sample of participants. Future studies could include a larger and more diverse participant pool to examine whether the patterns we observed hold across different levels of expertise or reading strategies. It is also worth exploring the long-term development of critical thinking skills following the study, without interface support, to better capture reading and critique behaviors without explicit scaffolding.
Third, LLM-based agents may generate hallucinated statements or adopt incorrect disciplinary framings. To mitigate this risk, we designed a reference feature for each agent-generated statement to encourage readers to verify claims against the source paper. However, it is inevitable that hallucinations cannot be fully prevented in LLMs. Rather than trying to eliminate hallucinations entirely, future work could explore interaction designs that help users rely on AI outputs more appropriately and think more critically about agent responses.
@@ Acknowledgements.
We thank Zefan Sramek, Shitao Fang, Jo Takezawa, and the other members of IIS Lab for their support during this work. This work was partly supported by JST ASPIRE for Top Scientists (Grant Number JPMJAP2405), JST PRESTO (Grant Number JPMJPR23IB), JST SPRING (Grant Number JPMJSP2108) and the JSPS Invitational Fellowship for Research in Japan program (Long-term, Grant Number L24509).
@@ Appendix D Prompts
Note that none, single, and multi condition share the same prompt for critical thinking questions generation. In single condition, the prompt to generate single-disciplinary feedback differs slightly from that in multi condition: it provides feedback from one disciplinary perspective that is different from the participant’s major, and without {pinned agent} input. The text inside {} represents the participants’ input.
D.1. Prompt for Critical Thinking Questions Generation
### **Input**
Section: {text}
Category: {category}
### **Output**
- Question
### **Requirement**
1. Read the given section content carefully.
2. Identify the main topic and the core ideas.
3. Generate **only one** critical thinking question that is directly relevant to the section and aligned with the assigned category.
4. Ensure that the question follows the structure and tone of the example questions provided.
5. Return **only the question** with no extra formatting or explanation.
6. Ensure the question is:
- Clear and easy to understand for junior researchers, using simple and precise language.
- Logically structured and specific, avoiding vague or overly abstract expressions.
- Thought-provoking and rigorous, encouraging the reader to reflect critically.
### **Examples**
**Input**
Section: {text}
Category: {category}
**Output**
{few_shot_questions}
D.2. Prompt for Multi-Disciplinary Agent Feedback
The instructor asks a critical thinking question ’{question}’, and the readers answer is ’{answer}’.
Judge the readers answer first, then provide an improved response from **experts in three distinct and relevant academic disciplines**.
First generate answers from ’{concatenated_disciplines}’s perspectives.
Each answer should emphasise how it **aligns with or challenges current trends or theories** in that discipline.
Ensure that each disciplines answer introduces fundamentally different insights rather than paraphrasing the same idea in different terminology.
Each answer should:
- **Begin with an implicit evaluation** of the readers answer before transitioning into the improved response.
- **Highlight alignment with or challenges to current trends or theories** in that discipline.
- **Ensure each disciplines answer introduces fundamentally different insights**, avoiding rephrasing the same idea.
- **Be concise (less than 100 words per answer).**
In the output, replace any instance of The readers answer with Your answer’.
Provide the output as **raw JSON text**, without any surrounding Markdown formatting like triple backticks.
Do not split or abbreviate any disciplines. Each discipline in ’{concatenated_disciplines}’ should be treated as a distinct entity.
If fewer than three disciplines are provided, choose an additional relevant discipline to complete the set of three.
[Perspective Name]: [Judgement of answer][Improved answer],
[Perspective Name]: [Judgement of answer][Improved answer],
[Perspective Name]: [Judgement of answer][Improved answer]
D.3. Prompt for Reinterpreting Highlights via Multi-Disciplinary Agents
The reader highlights ’{quote}’ in the paper and marks it as ’{label}’.
Reinterpret ’{label}’ of readers highlight from ** 3 different academic discipliness perspective**.
First generate reinterpretation from ’{concatenated_disciplines}’s perspectives.
Each reinterpretation should emphasise how it **aligns with or challenges current trends or theories** in that discipline.
Ensure that each disciplines reinterpretation introduces fundamentally different insights rather than paraphrasing the same idea in different terminology.
Each reinterpretation should:
- **Highlight alignment with or challenges to current trends or theories** in that discipline.
- **Ensure each disciplines reinterpretation introduces fundamentally different insights**, avoiding rephrasing the same idea.
- **Be concise (less than 100 words per reinterpretation).**
Provide the output as **raw JSON text**, without any surrounding Markdown formatting like triple backticks.
Do not split or abbreviate any disciplines. Each discipline in ’{concatenated_disciplines}’ should be treated as a distinct entity.
If fewer than three disciplines are provided, choose an additional relevant discipline to complete the set of three.
[Perspective Name]: [Reinterpretation],
[Perspective Name]: [Reinterpretation],
[Perspective Name]: [Reinterpretation]
D.4. Prompt for Answering Comments via Multi-Disciplinary Agents
The reader highlights ’{quote}’ in the paper and adds comment: ’{comment}’.
If the reader asks a question, answer the question based on highlights from ** 3 most relevant different academic discipliness perspective**.
If the reader leaves a statement, discuss around readers highlight and comment from ** 3 most relevant different academic discipliness perspective**.
First generate a response from ’{pinned_agent}’s perspectives.
Each response should emphasize how it **aligns with or challenges current trends or theories** in that discipline.
Ensure that each disciplines response introduces fundamentally different insights rather than paraphrasing the same idea in different terminology.
Each response should:
- **Highlight alignment with or challenges to current trends or theories** in that discipline.
- **Ensure each disciplines response introduces fundamentally different insights**, avoiding rephrasing the same idea.
- **Be concise (less than 100 words per response).**
Provide the output as **raw JSON text**, without any surrounding Markdown formatting like triple backticks.
Do not split or abbreviate any disciplines. Each discipline in ’{pinned_agent}’ should be treated as a distinct entity.
If fewer than three disciplines are provided, choose an additional relevant discipline to complete the set of three.
[Perspective Name]: [Response],
[Perspective Name]: [Response],
[Perspective Name]: [Response]
D.5. Prompt for Revised Answer Feedback
Users initial answer is ’{initialAnswer}’, the revised answer is ’{revisedAnswer}’.
The feedback for the initial answer is ’{agreeText}’.
@@ D.5. Prompt for Revised Answer Feedback
Based on this feedback, evaluate how well the revised answer addresses the issues or suggestions mentioned.
Provide a concise and constructive feedback on the revised answer in 100 words.